|
ID |
Date |
Author |
Subject |
Text |
 |
|
120
|
Wed Oct 13 12:06:09 2021 |
RSS TD | Flanges cleaned | RSS & TD had some success cleaning the COF600
flange surfaces of the staining and residues
observed when the downstream COF600 section |
12x |
|
119
|
Wed Oct 13 11:59:23 2021 |
TD | [How to] Open COF600 copper wire seals |
Use the M10 jacking points provided top/bottom
and
|
 |
|
118
|
Tue Oct 12 13:54:21 2021 |
RSS TD + OD DR | CARME opened | With the assistance of Oemer and Davide all
of the M22 bolts removed from the downstream
section of the CARME chamber. To separate |
20x |
|
117
|
Tue Oct 12 09:22:37 2021 |
RSS TD + OG DR | Motors problem solved | Neg and ion cables were taken off and were
put in the box.
Top support was taken off. |
 |
|
116
|
Mon Oct 11 11:24:16 2021 |
RSS TD ML | Valve support and cable labeling | Split funnels were given to Oemer for cleaning.
Have received them back after lunch.
Valve support is installed again |
  |
|
115
|
Sat Oct 9 09:47:33 2021 |
CB | To do list for next week | 0. Give split-funnels to OG or Grossmontage
to clean. Mount back foot support for gate
valves. Critical when we split. |
|
|
114
|
Sat Oct 9 09:44:36 2021 |
CB RSS | Heaters removed + motors tests | Removed thermocouples, heaters, RE72 modules
and leftover baking tent. Support octagon
removed except topmost part which requires |
|
|
113
|
Fri Oct 8 17:56:35 2021 |
CB RSS | Motors test | Attempted first motor test in XHV. Wired
and connected limit switches and string potentiometer.
Set up raspberry pi to control |
 |
|
112
|
Thu Oct 7 08:40:09 2021 |
CB RSS | CARME in XHV | CARME reached 9E-12 this morning
Ion pumps on
|
|
|
111
|
Wed Oct 6 15:36:14 2021 |
CB RSS | Neg activation tests -3 | Attempted to activate 2xUHV1400 connected
to bottom C4, channel #2 using P 200 W, t
rise 30 min, t hold 90 min.
|
|
|
110
|
Wed Oct 6 09:47:30 2021 |
CB RSS | Neg activation tests -2 | Attempted activating 2xUHV1400 connected to
top C4, channel #3 with P 300 W (450 C?),
t rise 30 min, t hold 90
|
|
|
109
|
Wed Oct 6 09:39:59 2021 |
PB CB | Original dirty assembly of CARME internals | PB found pictures of the original dirty assembly,
before screws and components were shipped
to DL.
|
7x |
|
108
|
Tue Oct 5 17:30:01 2021 |
CB RSS | NEG activation tests | Found D2000 #1 activated and at 250C (25W)
to prevent getting dirty when other NEGs |
  |
|
107
|
Tue Oct 5 11:44:11 2021 |
CB, TD | SAES manuals & cable drawings | NEG power
SIP power
UHV1400 operating instructions |
8x |
|
106
|
Mon Oct 4 20:30:46 2021 |
RSS JM CB (remotely) | CARME baking status -21 |
Current temperature ~95 C (on manual
mode)
See plot and screen attached.
Current
pressure 2.86E-10
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
|
 |
|
105
|
Mon Oct 4 09:19:11 2021 |
JM RSS | Activation of NEG elements | 9:30
Ion pumps de-gassed by switching
on and off multiple times. Large pressure |
   |
|
104
|
Mon Oct 4 06:34:58 2021 |
RSS JM CB (remotely) | CARME baking status -20 |
Current temperature ~110 C (on manual
mode)
See plot and screen attached.
Current
pressure 3.20E-8
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
|
 |
|
103
|
Sun Oct 3 08:36:01 2021 |
RSS JM CB (remotely) | CARME baking status -19 |
08:36
Current temperature ~132 C (on
manual mode RE72-1: 132 C, RE72-2: 130 C)
See
plot and screen attached.
Current pressure |
 |
|
102
|
Sat Oct 2 22:14:14 2021 |
RSS JM CB (remotely) | CARME baking status -18 |
Current temperature ~132 C (on manual
mode RE72-1: 132 C, RE72-2: 130 C)
See plot
and screen attached.
Current pressure |
 |
|
101
|
Sat Oct 2 08:26:49 2021 |
RSS JM CB (remotely) | CARME baking status -17 |
Current temperature ~132 C (on manual
mode RE72-1: 132 C, RE72-2: 130 C)
See plot
and screen attached.
Current pressure |
 |